Harris Fabrics computes its plantwide predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 42,000 direct labor-hours would be required for the period’s estimated level of production. The company also estimated $509,000 of fixed manufacturing overhead cost for the coming period and variable manufacturing overhead of $3.00 per direct labor-hour. Harris’s actual manufacturing overhead cost for the year was $712,685 and its actual total direct labor was 42,500 hours.

Required:

Compute the company’s plantwide predetermined overhead rate for the year. (Round your answer to 2 decimal places.)
Answer #1

Calculation of plantwide predetermined overhead rate:

Predetermined overhead rate= Estimated overhead/estimated direct labor hours

Estimated overhead= (42000*3)+509000= 635000

Predetermined overhead rate= 635000/42000= $15.12

Predetermined overhead rate= $15.12